# Dark Transformer: A Video Transformer for Action Recognition in the Dark

### v1

@@ Abstract
Recognizing human actions in adverse lighting conditions presents significant challenges in computer vision, with wide-ranging applications in visual surveillance and nighttime driving. Existing methods tackle action recognition and dark enhancement separately, limiting the potential for end-to-end learning of spatiotemporal representations for video action classification. This paper introduces Dark Transformer, a novel video transformer-based approach for action recognition in low-light environments. Dark Transformer leverages spatiotemporal self-attention mechanisms in cross-domain settings to enhance cross-domain action recognition. By extending video transformers to learn cross-domain knowledge, Dark Transformer achieves state-of-the-art performance on benchmark action recognition datasets, including InFAR, XD145, and ARID. The proposed approach demonstrates significant promise in addressing the challenges of action recognition in adverse lighting conditions, offering practical implications for real-world applications.
@@ 1 Introduction
Recognizing human actions across diverse environments presents a challenging task in computer vision. Over the past decade, numerous automated action recognition approaches have been developed, employing various modalities [1, 2, 3]. Human action recognition in dark conditions is crucial for applications like airport screening or mask-wearing detection and surveillance, especially during global challenges like the pandemic [4], where accurate action recognition is essential for ensuring public health and safety. Traditional feature-based and machine-learning methods have exhibited limited performance within specific domains, leading researchers to focus on deep learning-based techniques [5]. These techniques enable end-to-end processing of data for simultaneous feature extraction and classification. Convolutional Neural Network (CNN)-based methods have emerged as a prominent approach, leveraging hierarchical feature extraction. Initial layers capture local features, while subsequent layers extract global features. This paper provides an overview of the advancements in deep learning-based action recognition techniques, emphasizing the hierarchical feature extraction capabilities of CNNs.
@@
Despite their remarkable success in various application scenarios, the generalisation performance of deep neural networks often suffers when encountering new domains due to the domain shift problem. Unsupervised Domain Adaptation (UDA) aims to overcome this challenge by transferring knowledge from a labelled source domain to an unlabeled target domain. As documented in [7, 8, 9, 10, 11], extensive research efforts are dedicated to addressing the Unsupervised Domain Adaptation problem and alleviating the need for laborious annotations. These endeavours strive to enhance the adaptability and robustness of deep neural networks across diverse domains.
@@
The Transformer architecture is a suitable model because it explicitly builds contextual support for its representations through self-attention. This architecture has been tremendously effective for sequence modelling tasks compared to traditional recurrent models. It has also achieved remarkable success with visual tasks, and various architectures are being devised for image classification, segmentation, object detection, and video classification. Similarly, cross-domain transformers have been devised for image classification. However, there is still a research gap about cross-domain video transformers.
@@ 1.1 Related Work
Alternatively, another approach is to fuse multiple spectrums to address this issue. Previous works [18, 19, 2] have proposed methods for simultaneous action recognition from multiple video streams. Recently, Xu et al. [20] introduced the ARID dataset, which focuses on human actions in dark videos and reveals the ineffectiveness of current action recognition models and frame enhancement techniques in dark conditions. Dark-Light networks [21] leverage dark videos and their brightened counterparts, employing a dual-pathway structure for improved video representation and significantly enhancing action recognition performance. However, this approach is computationally intensive.
@@ 2 The Prpposed Method
Subsequently, in the subsequent subsections, we delve into the process of video tokenization and introduce the space-time cross-attention module. We explain how the input videos are tokenized into smaller units to enable efficient processing and information extraction. Additionally, we detail the implementation of the space-time cross-attention module, which facilitates the model’s ability to capture both spatial and temporal dependencies in an interconnected manner.
Discussing these aspects, we aim to thoroughly explain the model’s design choices and highlight the significance of video tokenization and space-time cross-attention in achieving robust action recognition performance.
@@ 2.1 Overview of Dark Transformer
The second branch, the "Target" branch, handles the target domain videos. This branch specifically deals with the challenges the target domain poses, such as adverse lighting conditions, and adapts the model to recognize actions effectively in such environments. The Target branch utilizes the learned knowledge from the Source branch while incorporating domain-specific transformations and adjustments to enhance its performance in the target domain.
Lastly, the "Source-Target" branch integrates information from both the Source and Target branches, facilitating learning domain-invariant representations. This branch aims to align and fuse the learned knowledge from both domains, enabling the model to generalize well across domains and exhibit robustness in action recognition tasks.
The framework’s inputs are selected pairs of related video sequences., $X_{1},X_{2}$, where $X_{1}\in D_{source}$, with a given labelled source (daytime) dataset $D_{1}$ and $X_{2}\in D_{target}$, an unlabelled target dataset (nighttime) $D_{2}$, with two different data distributions due to domain shift but they share an identical label space. We aim to train a model that can utilize the valuable source domain knowledge $D_{source}$ to help learn in the target domain $D_{target}$. Each video $X\in R^{T}\times H\times W\times N$ where N is the number of frames.
As shown in Fig. 3, the source and target videos in the input pair are sent to the source and target branches, respectively. In these two branches, the space-time self-attention module is involved in learning the domain-specific representations. The softmax cross-entropy loss is used to train the action classification. It is worth noting that all three branches share the same classifier due to the same label of two videos.
Figure 2: Architectural diagram of the proposed Dark Transformer Transformer during training. The model inputs source (daytime) and target domain (nighttime) videos. However, during inference, only the target domain videos are utilized as a single modality for action recognition.
@@ 2.2 Embedding Video Clips
We employ two straightforward methods to map a video, denoted as X, into a sequence of tokens, denoted as Z. The first method involves tokenizing the input video by uniformly sampling frames from the video clip. Each frame is then independently embedded into 2D tokens using the same method as ViT (Vision Transformer). These tokens are concatenated together to form the sequence representation. To handle each frame, we decompose it into M non-overlapping patches, where each patch has a size of S x S. These patches are flattened into vectors, denoted as v(s, t), where s ranges from 1 to M and t ranges from 1 to N, representing the spatial and temporal locations, respectively. Next, each patch is linearly embedded into embedding tokens, denoted as $Z(s,t)$, with additional spatiotemporal position embeddings. These embedding tokens are then passed through the transformer encoder for further processing.
@@
where the projection by E is equivalent to a 2D convolution, and P is learned positional embedding.
@@ 2.3 Space-Time self attentions (S+T)
This model forwards all Spatio-temporal tokens $Z(s,t)$ extracted from the video through the Dark Transformer transformer encoder. We calculate spatiotemporal self-attention for individual domains for source and target branches separately. The proposed model consists of $L$ encoding blocks. At each block, a query/key/value vector is computed for each patch from the representation $(s;t)$ encoded by the preceding block as follows:
@@
Where h is an index over multiple attention heads, and D is the latent dimension. As the direct computation of spatiotemporal self-attention is very expensive, we used a similar strategy as [] of using divided space and time attentions (S+T), where temporal and spatial attention are separately applied one after the other. This method computes temporal attention by comparing each patch with all the patches at the same spatial location in the other frames. Instead of being passed to the MLP, the encoding resulting from applying temporal attention is fed back for spatial attention computation.
@@ 2.4 Space-Time Cross attentions
Figure 3: The architectural diagram showcases the three weight-sharing branches of the Dark Transformer Transformer. The middle branch, highlighted in this diagram, plays a crucial role by providing accurate alignment through space-time cross-domain attention.
The cross-attention module makes the source-target branch features not only align distributions of two domains but also robust to noise in the input pairs. Thus, we utilize the output of the source-target branch as a training guide for the target branch.
@@ 3 Experimental Results
textbfDatasets. This study utilizes two datasets in the visible-to-infrared action recognition task: the InFAR dataset and the XD145 dataset. The XD145 dataset, which will be made available at https://sites.google.com/site/yangliuxdu/, serves as the source domain, while the InFAR dataset is employed as the target domain. These datasets provide valuable resources for training and evaluating the proposed approach, allowing for a comprehensive analysis of the model’s performance in cross-domain action recognition tasks from visible to infrared domains.
@@
(B) XD145 A visible light action dataset named XD145 reflects the above inFar dataset. In correspondence with the target domain action categories, the XD145 and the InfAR dataset have the same action categories, as shown in Figure 4. The XD145 action dataset consists of 600 video sequences captured by visible light cameras, and there are 50 video clips for each action class. All actions were performed by 30 different volunteers. Each clip lasts for 5 s on average. The frame rate is 25 fps, and the resolution is 320 × 240. As shown in Figure 4, the background, pose, and viewpoint variations are considered when constructing the dataset to make our dataset more representative of real-world scenarios.
@@
We follow a standard procedure during the inference phase to process the video clips from the target domain. This involves applying our trained network to the target domain video, analyzing each frame or video clip individually, and aggregating the results to obtain the final prediction or output. This procedure ensures that the network can adapt and generalize well to unseen videos in the target domain, enhancing our approach’s overall performance and accuracy in action recognition tasks.
@@
Similarly, for the InfAR dataset, Dark Transformer demonstrates a remarkable improvement of 6.83 % in performance compared to the 3D-ResNet-101 architecture. This notable enhancement further emphasizes the effectiveness of our approach in tackling action recognition in adverse lighting conditions.
@@ 3.1 Ablation study
[…]
We try to visualize feature visualization to illustrate further the performance of the domain-invariant nature of learned features. Fig 8 illustrates that our model’s feature visualization remains almost invariant for both domains.
Figure 7: Attention visualisation shows the domain-in-variance nature of learning by the proposed model as it shows the similarity between activation.
@@
3.2 Blind review
Many authors misunderstand the concept of anonymizing for blind review. Blind review does not mean that one must remove citations to one’s work—in fact, it is often impossible to review a paper unless the previous citations are known and available.
@@ 4 Conclusions
In conclusion, this paper introduces a novel approach for recognizing human actions in challenging lighting conditions by leveraging a video transformer-based model. The proposed model, named Dark Transformer, utilizes spatiotemporal self-attention in cross-domain settings, enabling the learning of cross-domain action recognition. By extending the capabilities of video transformers to incorporate cross-domain knowledge, Dark Transformer achieves state-of-the-art performance on multiple benchmark action recognition datasets.
Extensive experimentation and evaluation demonstrate Dark Transformer’s superior performance on the InFAR, XD145, and ARID datasets, renowned for their challenging lighting conditions and diverse actions. By effectively leveraging spatiotemporal self-attention and incorporating cross-domain knowledge, Dark Transformer surpasses previous approaches and sets a new benchmark for action recognition in adverse lighting conditions.
The results highlight the efficacy and potential of video transformer-based methods in tackling the unique challenges posed by varying lighting conditions, paving the way for future advancements in action recognition research and applications.

### v2

@@ Abstract
Recognising human actions in adverse lighting conditions presents significant challenges in computer vision, with wide-ranging applications in visual surveillance and nighttime driving. Existing methods tackle action recognition and dark enhancement separately, limiting the potential for end-to-end learning of spatiotemporal representations for video action classification. This paper introduces Dark Transformer, a novel video transformer-based approach for action recognition in low-light environments. Dark Transformer leverages spatiotemporal self-attention mechanisms in cross-domain settings to enhance cross-domain action recognition. By extending video transformers to learn cross-domain knowledge, Dark Transformer achieves state-of-the-art performance on benchmark action recognition datasets, including InFAR, XD145, and ARID. The proposed approach demonstrates significant promise in addressing the challenges of action recognition in adverse lighting conditions, offering practical implications for real-world applications.
@@ 1 Introduction
Recognising human actions across diverse environments presents a challenging task in computer vision. Over the past decade, numerous automated action recognition approaches have been developed, employing various modalities [1, 38, 3, 35]. Human action recognition in dark conditions is crucial for applications like airport screening or mask-wearing detection and surveillance, especially during global challenges like the pandemic [4], where accurate action recognition is essential for ensuring public health and safety. Traditional feature-based and machine-learning methods have exhibited limited performance within specific domains, leading researchers to focus on deep learning-based techniques [5]. These techniques enable end-to-end processing of data for simultaneous feature extraction and classification. Convolutional Neural Network (CNN)-based methods have emerged as a prominent approach, leveraging hierarchical feature extraction. Initial layers capture local features, while subsequent layers extract global features. This paper provides an overview of the advancements in deep learning-based action recognition techniques, emphasising the hierarchical feature extraction capabilities of CNNs.
@@
Despite their remarkable success in various application scenarios, the generalisation performance of deep neural networks often suffers when encountering new domains due to the domain shift problem. Unsupervised Domain Adaptation (UDA) aims to overcome this challenge by transferring knowledge from a labelled source domain to an unlabeled target domain. As documented in [7, 8, 9, 10, 11, 33], extensive research efforts are dedicated to addressing the Unsupervised Domain Adaptation problem and alleviating the need for laborious annotations. These endeavours strive to enhance the adaptability and robustness of deep neural networks across diverse domains.
@@
The Transformer architecture is a suitable model because it explicitly builds contextual support for its representations through self-attention. This architecture has been tremendously effective for sequence modelling tasks compared to traditional recurrent models. It has also achieved remarkable success with visual tasks, and various architectures are being devised for image classification, segmentation, object detection, and video classification. Similarly, cross-domain transformers have been devised for image classification. However, there is still a research gap about cross-domain video transformers [39].
@@ 1.1 Related Work
Alternatively, another approach is to fuse multiple spectrums to address this issue. Previous works [18, 19, 38] have proposed methods for simultaneous action recognition from multiple video streams. Recently, Xu et al. [20] introduced the ARID dataset, which focuses on human actions in dark videos and reveals the ineffectiveness of current action recognition models and frame enhancement techniques in dark conditions. Dark-Light networks [21, 38] leverage dark videos and their brightened counterparts, employing a dual-pathway structure for improved video representation and significantly enhancing action recognition performance. However, this approach is computationally intensive.
@@ 2 The Prpposed Method
Subsequently, in the subsequent subsections, we delve into the process of video tokenisation and introduce the space-time cross-attention module. We explain how the input videos are tokenised into smaller units to enable efficient processing and information extraction. Additionally, we detail the implementation of the space-time cross-attention module, which facilitates the model’s ability to capture both spatial and temporal dependencies in an interconnected manner.
Discussing these aspects, we aim to thoroughly explain the model’s design choices and highlight the significance of video tokenisation and space-time cross-attention in achieving robust action recognition performance.
@@ 2.1 Overview of Dark Transformer
The second branch, the "Target" branch, handles the target domain videos. This branch specifically deals with the challenges the target domain poses, such as adverse lighting conditions, and adapts the model to recognise actions effectively in such environments. The Target branch utilises the learned knowledge from the Source branch while incorporating domain-specific transformations and adjustments to enhance its performance in the target domain.
Lastly, the "Source-Target" branch integrates information from both the Source and Target branches, facilitating learning domain-invariant representations. This branch aims to align and fuse the learned knowledge from both domains, enabling the model to generalise well across domains and exhibit robustness in action recognition tasks.
The framework’s inputs are selected pairs of related video sequences., $X_{1},X_{2}$, where $X_{1}\in D_{source}$, with a given labelled source (daytime) dataset $D_{1}$ and $X_{2}\in D_{target}$, an unlabelled target dataset (nighttime) $D_{2}$, with two different data distributions due to domain shift but they share an identical label space. We aim to train a model that can utilise the valuable source domain knowledge $D_{source}$ to help learn in the target domain $D_{target}$. Each video $X\in R^{T}\times H\times W\times N$, where N is the number of frames.
As shown in Fig. 3, the source and target videos in the input pair are sent to the source and target branches, respectively. In these two branches, the space-time self-attention module is involved in learning the domain-specific representations. The softmax cross-entropy loss is used to train the action classification. It is worth noting that all three branches share the same classifier due to the same label for two videos.
Figure 2: Architectural diagram of the proposed Dark Transformer Transformer during training. The model inputs source (daytime) and target domain (nighttime) videos. However, during inference, only the target domain videos are utilised as a single modality for action recognition.
@@ 2.2 Embedding Video Clips
We employ two straightforward methods to map a video, denoted as X, into a sequence of tokens, denoted as Z. The first method involves tokenising the input video by uniformly sampling frames from the video clip. Each frame is then independently embedded into 2D tokens using the same method as ViT (Vision Transformer). These tokens are concatenated together to form the sequence representation. To handle each frame, we decompose it into M non-overlapping patches, where each patch has a size of S x S. These patches are flattened into vectors, denoted as v(s, t), where s ranges from 1 to M and t ranges from 1 to N, representing the spatial and temporal locations, respectively. Next, each patch is linearly embedded into embedding tokens, denoted as $Z(s,t)$, with additional spatiotemporal position embeddings. These embedding tokens are then passed through the transformer encoder for further processing.
@@
where the projection by E is equivalent to a 2D convolution, and P is a learned positional embedding.
@@ 2.3 Space-Time self attentions (S+T)
This model forwards all spatiotemporal tokens $Z(s,t)$ extracted from the video through the Dark Transformer transformer encoder. We calculate spatiotemporal self-attention for individual domains for source and target branches separately. The proposed model consists of $L$ encoding blocks. At each block, a query/key/value vector is computed for each patch from the representation $(s;t)$ encoded by the preceding block as follows:
@@
Where h is an index over multiple attention heads, and D is the latent dimension. As the direct computation of spatiotemporal self-attention is very expensive, we used a similar strategy to using divided space and time attentions (S+T), where temporal and spatial attention are separately applied one after the other. This method computes temporal attention by comparing each patch with all the patches at the same spatial location in the other frames. Instead of being passed to the MLP, the encoding resulting from applying temporal attention is fed back for spatial attention computation.
@@ 2.4 Space-Time Cross attentions
Figure 3: The architectural diagram showcases the three weight-sharing branches of the Dark Transformer. The middle branch, highlighted in this diagram, plays a crucial role by providing accurate alignment through space-time cross-domain attention.
The cross-attention module makes the source-target branch features not only align distributions of two domains but also robust to noise in the input pairs. Thus, we utilise the output of the source-target branch as a training guide for the target branch.
@@ 3 Experimental Results
textbfDatasets. This study utilises two datasets in the visible-to-infrared action recognition task: the InFAR dataset and the XD145 dataset. The XD145 dataset, which will be made available at https://sites.google.com/site/yangliuxdu/, serves as the source domain, while the InFAR dataset is employed as the target domain. These datasets provide valuable resources for training and evaluating the proposed approach, allowing for a comprehensive analysis of the model’s performance in cross-domain action recognition tasks from visible to infrared domains.
@@
(B) XD145 A visible light action dataset named XD145 reflects the above inFar dataset. In correspondence with the target domain action categories, the XD145 and the InfAR dataset have the same action categories, as shown in Figure 4. The XD145 action dataset consists of 600 video sequences captured by visible light cameras, and there are 50 video clips for each action class. All actions were performed by 30 different volunteers. Each clip lasts for 5 seconds on average. The frame rate is 25 fps, and the resolution is 320 × 240. As shown in Figure 4, the background, pose, and viewpoint variations are considered when constructing the dataset to make our dataset more representative of real-world scenarios.
@@
We follow a standard procedure during the inference phase to process the video clips from the target domain. This involves applying our trained network to the target domain video, analysing each frame or video clip individually, and aggregating the results to obtain the final prediction or output. This procedure ensures that the network can adapt and generalise well to unseen videos in the target domain, enhancing our approach’s overall performance and accuracy in action recognition tasks.
@@
Similarly, for the InfAR dataset, Dark Transformer demonstrates a remarkable improvement of 6.83 % in performance compared to the 3D-ResNet-101 architecture. This notable enhancement further emphasises the effectiveness of our approach in tackling action recognition in adverse lighting conditions.
@@ 3.1 Ablation study
Varying the number of input frames In our previous experiments, we fixed the number of input frames at 32. Now, we increase and decrease the number of frames input to the model, thereby proportionally increasing the number of tokens. We observed a slight increase in performance (0.5 $\%$) for an increase in the number of frames to 64. However, a decrease in the number of frames adversely impacts the performance (5 $\%$) if the number of frames is reduced to 16.
Feature visualisation illustrates the performance of the domain-invariant nature of learned features. Figure 8 illustrates that our model’s feature visualisation remains almost invariant for both domains.
Figure 7: Attention visualisation shows the domain-in-variance nature of learning by the proposed model, as it shows the similarity between activations.
@@ 4 Conclusions
In conclusion, this paper presents Dark Transformer, a novel video transformer-based model designed to recognise human actions in challenging lighting conditions. By leveraging spatiotemporal self-attention in cross-domain settings, the model learns robust representations that support cross-domain action recognition. Dark Transformer extends the capability of video transformers by incorporating cross-domain knowledge and achieves state-of-the-art results on benchmark datasets, including InFAR, XD145 and ARID, which are known for adverse lighting and diverse action categories. Extensive experiments show that Dark Transformer consistently outperforms existing approaches and establishes a new benchmark for action recognition in low-light environments. These findings highlight the effectiveness of video transformer-based methods in handling illumination variability and point toward promising directions for future research and real-world deployment in visual surveillance and autonomous systems.